# Textual Knowledge Matters: Cross-Modality Co-Teaching

### v1

@@ 4.4 Hyper-parameters Analysis
Figure 5: Impact of hyper-parameters.
5 Limitations
Dependency on the Visual Lexicon. The effectiveness of our method largely depends on a rich visual lexicon. Although massive category names have been gathered from existing datasets, it might still lack the categories in specific tasks, such as the Herbarium 19 dataset Tan et al. (2019) designed for botanical research. In such a context, our method may fail to produce informative textual descriptions and thus produce inferior results.
Inheriting Flaws from Foundation Models. Owing to our reliance on LLMs and VLMs for constructing category descriptive texts, our method inevitably inherits biases and drawbacks from these foundational models. For example, as shown in Tab. 10, the CLIP model exhibits suboptimal performance (Zero-Shot) on the FGVC Aircraft dataset Maji et al. (2013). As a result, our method fails to generate informative text descriptions and cannot obtain improvement compared to the visual-only competitor SimGCD.
Methods All Old New
Zero Shot 15.9 10.8 20.9
GCD 45.0 41.1 46.9
SimGCD 54.2 59.1 51.8
Ours 50.8 44.9 53.8
@@ 6 Conclusion
Appendix
@@ A. Datasets Settings
In the main paper, we follow the GCD protocol Vaze et al. (2022) to determine the number of known classes in the dataset and the number of labeled data samples selected from these classes. As illustrated in Fig. 6, the blue segment denotes the labeled data, whereas the yellow segment encompasses the unlabeled data, which includes both known and unknown categories. We conduct extensive experiments across four generic image classification datasets, namely, CIFAR-10, CIFAR-100 Krizhevsky et al. (2009), ImageNet-100, and ImageNet-1K Deng et al. (2009), and four fine-grained image classification datasets, specifically, CUB Wah et al. (2011), Stanford Cars Krause et al. (2013b), Oxford Pets Parkhi et al. (2012), and Flowers102 Nilsback & Zisserman (2008). Detailed dataset information is provided in Table 11.
@@
ImageNet-100 Deng et al. (2009) 31.9K 50 95.3K 100
ImageNet-1K Deng et al. (2009) 321K 500 960K 1000
CIFAR-10 Krizhevsky et al. (2009) 12.5K 5 37.5K 10
CIFAR-100 Krizhevsky et al. (2009) 20.0K 80 30.0K 100
CUB Wah et al. (2011) 1.5K 100 4.5K 200
Stanford Cars Krause et al. (2013b) 2.0K 98 6.1K 196
Oxford Pets Parkhi et al. (2012) 0.9K 19 2.7K 37
Flowers102 Nilsback & Zisserman (2008) 0.3K 51 0.8K 102
Figure 6: Configuration of the Dataset for the GCD Task.
@@ B.1. Main Results
CIFAR10 Krizhevsky et al. (2009) 98.2$\pm$0.3 98.0$\pm$0.1 98.6$\pm$0.5
CIFAR100 Krizhevsky et al. (2009) 85.7$\pm$0.6 86.3$\pm$0.6 84.6$\pm$0.9
ImageNet-100 Deng et al. (2009) 88.0$\pm$0.7 92.4$\pm$0.3 85.2$\pm$1.2
ImageNet-1K Deng et al. (2009) 64.8$\pm$0.1 77.8$\pm$0.1 58.3$\pm$0.1
CUB Wah et al. (2011) 76.6$\pm$1.0 80.6$\pm$0.8 74.7$\pm$1.1
Stanford Cars Krause et al. (2013b) 86.9$\pm$0.6 87.4$\pm$1.0 86.7$\pm$0.4
Oxford Pets Parkhi et al. (2012) 95.5$\pm$0.1 93.9$\pm$0.6 96.4$\pm$0.1
Flowers102 Nilsback & Zisserman (2008) 87.2$\pm$1.7 90.7$\pm$0.9 85.4$\pm$2.5
@@
B.2. Evaluation on Varying Known Knowledge
In the main paper, we adhere to the GCD protocol Vaze et al. (2022) for setting the number of known classes in the dataset and the number of labeled data samples selected from these classes. The number of selected known categories is denoted as $K_{L}=R_{k}*K$. The variable $R_{l}$ represents the proportion of data labeled from each known category. Therefore, $R_{k}$ and $R_{l}$ together determine the amount of labeled data. Following Vaze et al. (2022), we set $R_{k}$ to 0.5 for most datasets, except for CIFAR-100, where $R_{k}$ is set to 0.8. $R_{l}$ is maintained at 0.5 across all datasets. We compare SimGCD Wen et al. (2023) with our TextGCD on the CIFAR-100 and CUB datasets to assess the influence of hyper-parameters $R_{k}$ and $R_{l}$. To minimize the effects of randomness, each experiment is conducted three times, and the average All accuracy is reported.
@@ B.2. Evaluation on Varying Known Knowledge
B.3. Evaluation on Pseudo Label Forms
@@
B.4. Generating Text for Images Using BLIP
BLIP Li et al. (2022b) introduces an optimization objective for generative tasks, facilitating the generation of textual descriptions for images. We explore using BLIP Li et al. (2022b) (blip-large) and BLIP2 Li et al. (2023a) (blip2-opt-2.7b) to generate textual content for images. For equitable comparison, we employ prompts such as “The name of this object/bird is” for tag generation and “The feature of this object/bird is” for attribute generation. These elements—three tags and two attributes—are then amalgamated to construct descriptive texts for images. As demonstrated in Tab. 14, although using BLIP2 for text generation shows some improvement on the general dataset CIFAR-100, the methods employing BLIP and BLIP2 yield results significantly inferior to our retrieval-based method on the fine-grained dataset CUB. In particular, our method surpasses the BLIP2-based method by 35.7% in All accuracy. This underperformance is attributed to BLIP’s tendency to overlook detailed content, thus limiting its accuracy in fine-grained class discovery.
@@ B.4. Generating Text for Images Using BLIP
B.5. Visualization of Feature Distributions
To further evaluate the effectiveness of TextGCD, we utilize t-SNE for visualizing feature representations generated by the backbone network. We compare the trained backbones of GCD, SimGCD, and our TextGCD on the Oxford Pets Parkhi et al. (2012) dataset. As depicted in Fig. 9 (a), (b), and (c)—all utilizing ViT-B as the backbone—it is evident that TextGCD produces more discriminative feature representations. Moreover, Fig. 9 (d) further validates the effectiveness of TextGCD’s textual feature representations.
@@ B.5. Visualization of Feature Distributions
C. Discussion and Future Direction
Visual Lexicon Selection. There is overlap between the datasets used for evaluation and those used to build the Visual Lexicon. Specifically, ImageNet-1K is utilized in both contexts. Our goal is to collect as much object semantics as possible to construct a rich and diverse Visual Lexicon. Given its popularity and the extensive common visual object semantics it contains, which are also prevalent in other datasets, ImageNet-1K’s class names are included when constructing the Visual Lexicon.
Tailored Visual Lexicon. Since labeled data are available in the GCD task, they can be used to confirm the scope of classified images, especially for fine-grained classification. We create a tailored Visual Lexicon for the Flower102 dataset by employing an LLM (Llama-2-13b) model to filter out tags unrelated to flowers and plants. As shown in Tab. 15, the tailored Visual Lexicon leads to improved results. In our future work, we plan to construct unique Visual Lexicons for each dataset.
@@ C. Discussion and Future Direction
D. Illustration
D.1. Visual Lexicon
@@
which has short-haired, stocky cat which has large, round eyes which has broad head and cheeks which has short, dense fur which has short, thick legs which has short, thick tail
D.2. Categorical Descriptive Text
@@
Figure 10: Generated categorical descriptive text: representative examples in generic datasets.
@@
Figure 11: Generated categorical descriptive text: representative examples in fine-grained datasets.

### v2

@@ 5 Conclusion
Acknowledgement
We acknowledge the support of the MUR PNRR project iNEST-Interconnected Nord-Est Innovation Ecosystem (ECS00000043) funded by the NextGenerationEU. Also, this work was partially supported by the EU Horizon projects ELIAS (No. 101120237) and AI4Trust (No. 101070190).
Supplementary Material
[…]
@@ A. Datasets Settings
In the main paper, we follow the GCD protocol [40] to determine the number of known classes in the dataset and the number of labeled data samples selected from these classes. We conduct extensive experiments across four generic image classification datasets, namely, CIFAR-10, CIFAR-100 [18], ImageNet-100, and ImageNet-1K [8], and four fine-grained image classification datasets, specifically, CUB [41], Stanford Cars [15], Oxford Pets [31], and Flowers102 [28]. Detailed dataset information is provided in Table 10.
@@
ImageNet-100 [8] 31.9K 50 95.3K 100
ImageNet-1K [8] 321K 500 960K 1000
CIFAR-10 [18] 12.5K 5 37.5K 10
CIFAR-100 [18] 20.0K 80 30.0K 100
CUB [41] 1.5K 100 4.5K 200
Stanford Cars [15] 2.0K 98 6.1K 196
Oxford Pets [31] 0.9K 19 2.7K 37
Flowers102 [28] 0.3K 51 0.8K 102
@@ B.1. Main Results
CIFAR10 [18] 98.2$\pm$0.3 98.0$\pm$0.1 98.6$\pm$0.5
CIFAR100 [18] 85.7$\pm$0.6 86.3$\pm$0.6 84.6$\pm$0.9
ImageNet-100 [8] 88.0$\pm$0.7 92.4$\pm$0.3 85.2$\pm$1.2
ImageNet-1K [8] 64.8$\pm$0.1 77.8$\pm$0.1 58.3$\pm$0.1
CUB [41] 76.6$\pm$1.0 80.6$\pm$0.8 74.7$\pm$1.1
Stanford Cars [15] 86.9$\pm$0.6 87.4$\pm$1.0 86.7$\pm$0.4
Oxford Pets [31] 95.5$\pm$0.1 93.9$\pm$0.6 96.4$\pm$0.1
Flowers102 [28] 87.2$\pm$1.7 90.7$\pm$0.9 85.4$\pm$2.5
@@
B.2. Additional Baseline Results
In Tab. 12, we present a comparison between the results of two baselines and our TextGCD. The term “Zero-shot” refers to the zero-shot results obtained using CLIP [33] in the GCD setting, given the names of all categories (including new categories). “Frozen-text” refers to partially fine-tuning the image encoder only while using frozen text embeddings of the category labels as classifiers. This approach treats the problem as partially supervised, where only some classes are unannotated, but all class names are known. As shown in Tab. 12, our TextGCD significantly outperforms these two baselines.
Table 12: Comparison of results between two baselines with known text descriptors and our TextGCD.
Methods CIFAR100 CUB Stanford Cars
All Old New All Old New All Old New
Zero-shot 66.2 64.6 69.5 54.3 62.3 50.3 67.3 59.4 71.3
Frozen-text 79.2 81.5 74.6 62.6 80.2 53.9 74.9 77.9 73.4
TextGCD 85.7 86.3 84.6 76.6 80.6 74.7 86.9 87.4 86.7
B.3. Evaluation on Varying Known Knowledge
In the main paper, we adhere to the GCD protocol [40] for setting the number of known classes in the dataset and the number of labeled data samples selected from these classes. The number of selected known categories is denoted as $K_{L}=R_{k}*K$. The variable $R_{l}$ represents the proportion of data labeled from each known category. Therefore, $R_{k}$ and $R_{l}$ together determine the amount of labeled data. Following [40], we set $R_{k}$ to 0.5 for most datasets, except for CIFAR-100, where $R_{k}$ is set to 0.8. $R_{l}$ is maintained at 0.5 across all datasets. We compare SimGCD [43] with our TextGCD on the CIFAR-100 and CUB datasets to assess the influence of hyper-parameters $R_{k}$ and $R_{l}$. To minimize the effects of randomness, each experiment is conducted three times, and the average All accuracy is reported.
@@ B.3. Evaluation on Varying Known Knowledge
B.4. Additional Results with Unknown $K$
In the main paper, we follow the common practice [43] of assuming the total number of categories is known a priori. Here, we explore the scenario where such knowledge is unknown. Vaze et al. [40] provides an off-the-shelf method to estimate the class number in unlabeled data. We introduce CLIP’s visual feature into the off-the-shelf method. As shown in Tab. 13, we present the estimated number of categories and the error rates for CIFAR100, CUB, and Stanford Cars. With the estimation of the number of categories, we compared our results with the SimGCD algorithm on these three datasets. As shown in Tab. 13, our TextGCD surpasses SimGCD (CLIP backbone) by an average of 9.9% in terms of All accuracy across the three datasets.
Table 13: Comparison with SimGCD with unknown category numbers. The upper part of the table shows the estimated number of categories, while the lower part presents the accuracy comparison.
CIFAR100 CUB Stanford Cars
Ground truth $K$ 100 200 196
Estimation 102 193 212
Error 2% 3.5% 8.2%
All Old New All Old New All Old New
SimGCD(DINO) 80.6 81.0 79.1 62.0 67.1 59.3 48.6 64.8 40.7
SimGCD(CLIP) 82.3 83.5 80.1 60.2 74.5 52.9 70.4 75.0 67.3
Ours 86.4 87.2 84.7 74.8 80.3 72.0 81.5 77.8 83.2
B.5. Evaluation on Pseudo Label Forms
@@
B.6. Generating Text for Images Using BLIP
BLIP [22] introduces an optimization objective for generative tasks, facilitating the generation of textual descriptions for images. We explore using BLIP [22] (blip-large) and BLIP2 [21] (blip2-opt-2.7b) to generate textual content for images. For equitable comparison, we employ prompts such as “The name of this object/bird is” for tag generation and “The feature of this object/bird is” for attribute generation. These elements—three tags and two attributes—are then amalgamated to construct descriptive texts for images. As demonstrated in Tab. 15, although using BLIP2 for text generation shows some improvement on the general dataset CIFAR-100, the methods employing BLIP and BLIP2 yield results significantly inferior to our retrieval-based method on the fine-grained dataset CUB. In particular, our method surpasses the BLIP2-based method by 35.7% in All accuracy. This underperformance is attributed to BLIP’s tendency to overlook detailed content, thus limiting its accuracy in fine-grained class discovery.
@@ B.6. Generating Text for Images Using BLIP
B.7. Visualization of Feature Distributions
To further evaluate the effectiveness of TextGCD, we utilize t-SNE for visualizing feature representations generated by the backbone network. We compare the trained backbones of GCD, SimGCD, and our TextGCD on the Oxford Pets [31] dataset. As depicted in Fig. 8 (a), (b), and (c)—all utilizing ViT-B as the backbone—it is evident that TextGCD produces more discriminative feature representations. Moreover, Fig. 8 (d) further validates the effectiveness of TextGCD’s textual feature representations.
@@ B.7. Visualization of Feature Distributions
C. Limitations
Dependency on the Visual Lexicon. The effectiveness of our method largely depends on a rich visual lexicon. Although massive category names have been gathered from existing datasets, it might still lack the categories in specific tasks, such as the Herbarium 19 dataset [39] designed for botanical research. In such a context, our method may fail to produce informative textual descriptions and thus produce inferior results.
Inheriting Flaws from Foundation Models. Owing to our reliance on LLMs and VLMs for constructing category descriptive texts, our method inevitably inherits biases and drawbacks from these foundational models. For example, as shown in Tab. 16, the CLIP model exhibits suboptimal performance (Zero-Shot) on the FGVC Aircraft dataset [26]. As a result, our method fails to generate informative text descriptions and cannot obtain improvement compared to the visual-only competitor SimGCD.
Methods All Old New
Zero Shot 27.0 21.3 32.3
GCD 45.0 41.1 46.9
SimGCD 54.2 59.1 51.8
Ours 50.8 44.9 53.8
D. Discussion and Future Direction
D.1. Visual Lexicon Selection
There is overlap between the datasets used for evaluation and those used to build the Visual Lexicon. Specifically, ImageNet-1K is utilized in both contexts. Our goal is to collect as much object semantics as possible to construct a rich and diverse Visual Lexicon. Given its popularity and the extensive common visual object semantics it contains, which are also prevalent in other datasets, ImageNet-1K’s class names are included when constructing the Visual Lexicon.
D.2. Tailored Visual Lexicon
Since labeled data are available in the GCD task, they can be used to confirm the scope of classified images, especially for fine-grained classification. We create a tailored Visual Lexicon for the Flower102 dataset by employing an LLM (Llama-2-13b) model to filter out tags unrelated to flowers and plants. As shown in Tab. 17, the tailored Visual Lexicon leads to improved results. In our future work, we plan to construct unique Visual Lexicons for each dataset.
@@ D.2. Tailored Visual Lexicon
D.3. Extend to SSL Setting
Our method can be easily extended to a semi-supervised (SSL) setting. Specifically, in a semi-supervised setting, the number of categories and the category names can be considered known. By applying our method to SSL, we can create a Visual Lexicon with ground-truth class names along with attributes obtained by LLMs (such as GPT-3). Based on this tailored Visual Lexicon, our TextGCD can then be applied to the SSL setting.
D.4. Extend to Incremental Setting
For GCD models to be effectively applied in real-world scenarios, Zhang et al. proposed the Continuous Category Discovery (CCD) task [50], which not only requires the ability to recognize data from incrementally increasing unknown categories but also emphasizes the importance of not forgetting previously known categories. Our method can be applied to the CCD task, as each training stage in CCD is a variant of GCD. The key to this extension is adding a tailored module to prevent forgetting the categories seen in previous training stages, such as the static-dynamic distillation loss in GM [50] or the distillation loss in iCaRL [34]. It is essential to maintain the memory of previously learned categories at each training stage. This is the design plan for extending TextGCD to the incremental setting.
E. Illustration
E.1. Visual Lexicon
@@
which has short-haired, stocky cat which has large, round eyes which has broad head and cheeks which has short, dense fur which has short, thick legs which has short, thick tail which has thick, dense coat
Egyptian Mau cat
which has short-haired cat which has spotted or marbled coat which has large eyes which has slender body which has muscular hindquarters which has long, thin tail
E.2. Categorical Descriptive Text
@@
Figure 9: Generated categorical descriptive text in the RTG phase: representative examples in generic datasets.
@@
Figure 10: Generated categorical descriptive text in the RTG phase: representative examples in fine-grained datasets.